OpenStudy (anonymous):

The quadratic equation whose roots are 1/(1 + root 3) and 1/(1- root3) is ?

OpenStudy (callisto):

(x-α)(x-β) = 0 , where α and β are the roots of the quad. equation. \[[x- \frac{1}{(1 + \sqrt3)} ][ x - \frac{1}{(1- \sqrt3)}] =0\]expand and simplify.

OpenStudy (anonymous):

YUP GOT IT